(3D^2-2D^2+D-1)z=4e^(x+y).cos(x+y)


1
Expert's answer
2021-05-11T11:40:26-0400

Given equation is-

"(3D^2-2D^2+D-1)z=4e^{(x+y)}.cos(x+y)\\[9pt](D^2+D-1)z=4e^{(x+y)}.cos(x+y)"


Auxiliary equation is-


"m^2+m-1=0"


"m=\dfrac{-1\pm \sqrt{1-4}}{2}=\dfrac{-1\pm \sqrt{3}i}{2}"


Complimentary function is-


"CF=e^{-\frac{x}{2}}(c_1cos \dfrac{\sqrt{3}}{2}x+c_2sin\dfrac{\sqrt{3}}{2}x)"



Particular integral is-

"PI=\dfrac{4e^{x+y}cos(x+y)}{D^2+D-1}"


"=4e^{x+y}\dfrac{cos(x+y)}{(D+1)^2+(D+1)-1}\\[9pt]=4e^{x+y}\dfrac{cos(x+y)}{D^2+3D+1}
\\[9pt]=4e^{x+y}\dfrac{cos(x+y)}{-1+3D+1}\\[9pt]=4e^{x+y}\dfrac{cos(x+y)}{3D}\\[9pt]=-4e^{(x+y)}\times \dfrac{sin(x+y)}{3}"


Complete solution is-


y=CF+PI

"y = e^{-\frac{x}{2}}(c_1cos \dfrac{\sqrt{3}}{2}x+c_2sin\dfrac{\sqrt{3}}{2}x)-4e^{(x+y)}\times \dfrac{sin(x+y)}{3}"